A section of iron wire having a diameter of 0.030 inch is measured to have a resistance of


2500


Ω (at a temperature of 20°C).


a. Determine the cross


-sectional area in square meters and in square millimeters.


(Note: 1 inch = 2.54 cm = 25.4 mm)


b. Calculate the length of the wire in meters.

1
Expert's answer
2022-03-17T09:47:25-0400

We know that

"R=\frac{\rho l}{A}"

"l=\frac{RA}{\rho}"


"l=\frac{2500\times3.14\times(3.81\times10^-4)^2}{9.81\times10^{-8}}=11615.83m"

Area

"A=\frac{\rho l}{R}"


"A=\frac{9.71\times10^{-8}\times11615.83}{2500}=4.52\times10^{-7}m^2"

"A=0.452mm^2"